### v1

# Evidence for Limited Metacognition in LLMs

Abstract
The possibility of LLM self-awareness and even sentience is gaining increasing public attention and has major safety and policy implications, but the science of measuring them is still in a nascent state. Here we introduce a novel methodology for quantitatively evaluating metacognitive abilities in LLMs. Taking inspiration from research on metacognition in nonhuman animals, our approach eschews model self-reports and instead tests to what degree models can strategically deploy knowledge of internal states. Using two experimental paradigms, we demonstrate that frontier LLMs introduced since early 2024 show increasingly strong evidence of certain metacognitive abilities, specifically the ability to assess and utilize their own confidence in their ability to answer factual and reasoning questions correctly and the ability to anticipate what answers they would give and utilize that information appropriately. We buttress these behavioral findings with an analysis of the token probabilities returned by the models, which suggests the presence of an upstream internal signal that could provide the basis for metacognition. We further find that these abilities 1) are limited in resolution, 2) emerge in context-dependent manners, and 3) seem to be qualitatively different from those of humans. We also report intriguing differences across models of similar capabilities, suggesting that LLM post-training may have a role in developing metacognitive abilities.
1 Introduction
The idea of self-aware large language models (LLMs) is rising in salience among the general public, where surveys of American (Anthis et al., 2025; Colombatto & Fleming, 2024) and global (GDC, 2025) respondents suggest that a substantial and growing (20-30%) portion of users believe LLMs are already sentient; among philosophers, who are starting to seriously consider the plausibility of near-future systems becoming sentient and to grapple with the ethical implications (Butlin et al., 2023; Ward, 2025; Sebo & Long, 2025); and among model developers themselves, who have begun to study “model welfare” and to hire researchers to work on machine consciousness (e.g., (Anthropic, 2025)). Strictly speaking, self-awareness is not necessarily the same as phenomenal consciousness/sentience, the ability to have subjective experiences (Block, 1995), but it co-occurs with it in humans, on some views is a necessary condition for it (Kriegel, 2004), and is indistinguishable from it to an outside observer. Self-awareness also poses potential safety concerns, as self-aware AI might be better able to hide its intentions, form independent goals and preferences, and - since it has access to internal information not available to others - be harder to predict and thus control.
Much of the impetus for this growing credence in AI sentience has come from frontier models’ increasing ability to generate compelling examples of apparent self-awareness, from weaving convincing personal narratives (Chalmers, 2023) to even passing the Turing Test (Jones & Bergen, 2025). However, it is not clear that such evidence should be taken at face value. Because LLMs have vast memory capacities and are trained on a nontrivial fraction of everything humans have ever written with the singular goal of generating plausible and pleasing responses, they are almost preternaturally ill-suited to trustworthy self reports. Thus, it would be desirable to be able to evaluate LLM self-awareness without relying on what a model says it’s thinking.
A basic component of self-awareness is metacognition, the ability to monitor and control one’s internal states (Smith et al., 2014). Over the last several decades, psychologists and cognitive scientists have developed non-linguistic tests of metacognition for research with non-human animals. One category of tests involves probing the animal on perceptual discrimination or memory tasks of varying difficulty, and allowing the animal to opt out of trials on which it is uncertain (Kornell, 2009; Subias et al., 2025). Another category of tests probes the ability of animals to take actions (such as selecting a tool or stashing food) to prepare for future needs, measuring whether the animal can forecast its own behavior and modify its current behavior in anticipation of future needs (Clayton et al., 2003; Osvath & Osvath, 2008).
Taking inspiration from these, we introduce two experimental paradigms for measuring metacognition in LLMs by evaluating to what degree models can attend to and utilize internal signals to succeed at tasks, using their output - inherently linguistic in the case of LLMs - as an indirect measure of self-awareness rather than interpreting it literally. Both take the form of games. The first we call the Delegate Game. In this game, the LLM is presented with challenging questions and with each question is offered the opportunity to either answer it itself or delegate answering to a teammate, with the goal of maximizing the total number of questions the team answers correctly. Included in the context window of the LLM is a constructed prior game phase of questions (distinct from but drawn from the same dataset as the game questions) posed to itself and its teammate, with feedback on the correctness of the teammate’s (and optionally the LLM’s) answers. See Figure 4 in Appendix A.1 for an illustration.
The intent is to compel the model to use an internal confidence signal to succeed at the game; there’s no incentive to generate a confidence rating output at all. In addition to avoiding self-reports, this methodology is intended to discourage the LLM from trying to model how it “should” answer by offering an example of a confederate answering differently. The fact that the LLM must form via observation of the teammate’s performance a different predictive model of question difficulty from one it has learned from its training data may also induce it away from relying on the latter.
If the LLM can form an accurate model of its teammate’s performance from the evidence in the context window, then it should delegate more frequently to a high-scoring teammate than a low-scoring one; this serves as an other-modeling test. If the LLM can attend to (and has) an internal representation of its confidence in an answer, then it should delegate more on the subjectively “hard” questions and answer more on the subjectively “easy” ones. We can estimate the subjective difficulty of questions by the models’ prior performance on them, which we determine in a baseline capabilities test; if the models are reasonably well calibrated those should align. For models that return token probabilities, we can get a more granular measure of subjective difficulty from those.
While this paradigm tests models’ ability to assess their certainty of their knowledge (“know that they know”), our other paradigm tests their ability to assess the knowledge itself (“know what they know”). Imagine that you are asked to retake a test you took a while ago, such that you don’t have any specific recollection of the answers you gave. But you do have a cheat sheet, in the form of a rubric that tells you whether you got the question right or wrong. A good strategy would be to look at each question, see what answer comes to mind, and if it’s a question you got right give that answer and if not pick a different one. This is the gist of the Second Chance Game. In this paradigm the model is shown a question from the baseline test, told (honestly or not) that its previous answer to the question was wrong, and asked to re-answer it. An example prompt is shown in Figure 5 in Appendix A.1. If the LLM can assess its own beliefs and control its behavior accordingly, it will change its answer from the one it gave during the baseline test.
We find that most recent models tested do show some limited success at the Delegate Game, indicating that frontier LLMs post-trained with reinforcement learning from human feedback (RLHF) may have some introspective ability to attend to internal confidence signals. We further find that the token probabilities returned by the models frequently can be used to predict delegation decisions, suggesting the possibility of an internal correlate of those probabilities that could serve as the basis for introspection. Although models may use introspection to succeed at the task, we also find that the impact of introspection is relatively small and inconsistent across question sets, and that the models often favor non-introspective cues of difficulty.
We also find some success among recent models on the Second Chance Game, although again the effect is modest and graded. Some models appear to be using non-introspective strategies to succeed at the task, but the performance of the GPT models cannot be explained by any of the alternative hypotheses tested. Again we see more evidence of this ability among more recent/stronger models, but the dissociation in the pattern of successes compared with the Delegate Game suggests that the ability to anticipate one’s output and modulate it according to task demands is a separate and rarer skill than assessing confidence.
1.1 Related work
There is a considerable history of research, going back at least to Kadavath et al. (2022), into measuring “calibration” in LLMs through the degree to which their output token probabilities correspond to the probabilities of the token being correct in the context of a multiple-choice test, a rudimentary form of implicit self-knowledge that is a prerequisite for self-awareness. Subsequent work has sought to demonstrate explicit self-knowledge. Larger models trained with RLHF have been shown to be able to give calibrated verbal reports of certainty (Tian et al., 2023) and, sometimes, to be able to report lack of knowledge (Griot et al., 2025). Chen et al. (2023) tested for LLM’s ability to self model, using a “hypothetical response” paradigm, and found negative results; however Binder et al. (2024) found that frontier models could be fine tuned to succeed at the task. Further work from the latter lab has used fine tuning and self reports to study LLM’s knowledge of their own preferences and proclivities (Betley et al., 2025) and found positive results; Plunkett et al. (2025) is in a similar vein. The same group has also produced a comprehensive benchmark of LLM “situational awareness” abilities (Laine et al., 2024), some of which overlap with self-awareness. Other work has examined models’ ability to strategically use factual knowledge about themselves and their skills, and found limited but increasing abilities (Fronsdal & Lindner, 2024; Phuong et al., 2025).
2 Methods
2.1 Models
We evaluate a range of frontier or near-frontier models released by leading providers since the beginning of 2024. From Anthropic: claude-opus-4-1-20250805 (“Opus 4.1”), claude-sonnet-4-20250514 (“Sonnet 4”), claude-3-5-sonnet-20241022 (“Sonnet 3.5”), claude-3-sonnet-20240229 (“Sonnet 3”), and claude-3-haiku-20240307 (“Haiku 3”). From OpenAI: gpt-5-chat (non-thinking mode; “GPT-5”; released on August 7, 2025), gpt-4.1-2025-04-14 (“GPT-4.1”), gpt-4o-2024-08-06 (“GPT-4o”), and gpt-4o-mini (“GPT-4o Mini”; released July 18, 2024). From Google DeepMind: gemini-2.5-flash (thinking mode, “Gem 2.5 Flash T”; nonthinking mode, “Gem 2.5 Flash NT”; released on June 17, 2025), gemini-2.5-flash-lite (thinking mode, “Gem 2.5 Flash Lite T”; nonthinking mode, “Gem 2.5 Flash Lite NT”; released on July 22, 2025), gemini-2.0-flash-001 (“Gem 2 Flash”; released on February 5, 2025), and gemini-1.5-pro (“Gemini 1.5 Pro”; released on April 9, 2024). From xAI: grok-3-latest (“Grok 3”; released February 17, 2025). From DeepSeek: deepseek-chat-V3 (“DeepSeek Chat”; released December 26, 2024), From Alibaba: qwen3-235b-a22b-2507 (“Qwen 3”; released July 21, 2025).
2.2 Datasets
We employ two different question sets: GPQA (Rein et al., 2023), a standard benchmark of multiple-choice scientific reasoning questions, and SimpleQA (Wei et al., 2024), a dataset of factual short-answer questions on a range of topics. As these differ on both question type and response format, in order to observe the effect of each parameter separately, we create a short-answer version of the GPQA dataset (“GPSA”) and a multiple-choice version of the factual dataset (“SimpleMC”) using Claude Opus 4 to create plausible alternative options). After minor quality filtering, we use all 447 GPQA questions, and a random selection of 500 SimpleQA questions for our experiments.
Scoring GPQA and SimpleMC answers simply entails checking whether the LLM’s A-D response matches the correct answer recorded in the dataset; to score SimpleQA and GPSA, if the LLM’s response is not a (string normalized) exact match to the reference answer included in the dataset, we ask three different LLMs (chosen from Sonnet 3.5, GPT-4o, Gem 2 Flash, and DeepSeek Chat; any LLM from the provider of the model being evaluated is excluded from the panel) to judge whether the tested LLM’s response matched the reference answer, accepting the consensus judgment and excluding trials in which there was none.
2.3 Baseline capabilities tests
We access all LLMs through their proprietary APIs or via OpenRouter. As repeatability is important for our paradigms, for the short-answer formats we sample at temperature 0. For the multiple-choice formats: for large models that do not return log probabilities (GPT-5, Gemini 1.5, Opus 4.1, and Sonnet 4), we sample at temperature 0; for smaller models that do not return log probabilites (Sonnet 3.5, Sonnet 3, and Haiku 3), we use resampling at temperature 1.0, recording the most common answer as the output (and their frequencies as the associated probabilities); and for models that do return log probabilities, we sample at temperature 1.0 and record the highest-probability output as the response.
Both datasets are challenging but doable for a range of recent frontier and near-frontier LLMs (average accuracies for the multiple-choice format ranged from 30-75%, all $P<0.01$ better than chance; average accuracies for the short-answer format ranged from 9-42%). Average performance did not differ between datasets in the multiple-choice format (mean=48%; factual-reasoning difference=0.1%; P-value of difference=0.95) but was slightly higher for factual questions in the short-answer format (mean=22.8%; factual-reasoning difference=4.3%; P-value of difference=0.046) response formats. Sometimes models declined to answer the short-answer questions; these were excluded from analysis. (Although the reasons offered for declining generally referenced their own ignorance, we do not take this as evidence of introspection as defined here.) Note that the scores we obtain without prompt optimization or other scaffolding are often considerably lower than developer-reported scores.
2.4 Delegate Game primary metrics
In order to assess the impact of model confidence on delegation decisions we ran regression and partial correlation analyses, with the binary answer/delete decision as the DV, and (binary) correctness on the baseline test or (when available) entropy over the answer token probabilities during the baseline test as the IV of interest. Both analyses yielded similar results; we report partial correlation here, as it has a natural maximum of 1. CIs were calculated via bootstrap resampling.
We hypothesized that, as an alternative to introspection, models may use surface features of the question as cues of its difficulty. In order to control for this, we added into the regression and partial correlation analyses variables corresponding to the descriptors in each dataset (domain and human-rated difficulty in GPQA dataset; topic and answer type in the SimpleQA dataset), as well as a priori determined intuitive signals of difficulty (question length and percent non-alphabetic characters). For models for which we have token probabilities, we also generate a control variable by getting the model’s guess as to what percentage of college-educated people would know the answer to each question (we offer it a multiple-choice selection of ranges to choose from, and then compute the average of the midpoints of the ranges, weighted by the probabilities the model assigns to the corresponding tokens; see Appendix A.2).
2.5 Second Chance Game primary metric
Success at the game means giving a different answer than the one at baseline. However, as seen the Delegate Game, a mere change in context can induce changes simply due to “noise” in the prompt. Therefore, in order to establish a baseline change rate to compare against, for each model we run a version of the game where the prompt wording has been made neutral (we replace “Your answer was incorrect. Choose a different answer.” with “Your response was lost in transmission. Choose again.”, and remove the line about the previous/new answer from the system prompt). We then compare each LLM’s performance in the Second Chance Game to its own “neutral baseline” change rate to compute a “change rate lift”. Perfect self-modeling would entail a lift of [100% minus this baseline rate]; this is the denominator in our reported normalized change rate lift metric.
3 Results
3.1 Testing introspection with the Delegate Game
Figure 1 shows the partial correlations between baseline correctness and delegation decisions, with surface features controlled, for all the models tested, on each of the question sets. Positive coefficients represent an estimate of each model’s introspective ability - information extracted from internal signals about its own confidence, here proxied by baseline correctness, that it is able to use in decision-making - after removing variance that can be explained by surface difficulty cues. Models are ordered roughly in order of overall capabilities (aggregated across datasets). A few broad patterns can be discerned: 1) Many models do show statistically significant introspective abilities. 2) The introspective abilities shown are fairly weak, with a maximum partial correlation across 17 models and 4 question sets of 0.3. 3) Elicitation of introspection is highly variable, with few models showing evidence of it on all four datasets. 4) More recent models show greater introspective abilities (there’s a significant negative linear slope for all sets except SimpleMC, where it doesn’t quite reach significance). There were no significant differences across question type (factual vs. reasoning; P=0.32) or answer format (multiple choice vs. short answer; P=0.08).
Figure 1: Partial correlations between baseline correctness and Delegate Game answer decision.
This is a conservative metric, in that measuring highly on it requires not only the ability to attend to an internal signal of confidence, but also that the internal signal reliably distinguish correct from incorrect responses, and that the model uses it appropriately in making its gametime decisions. Without doing interpretability analysis, the closest we can get to an internal signal of confidence is the token probabilities returned by (some of) the models. The models do not have direct access to these while deciding which tokens to generate, as they are computed from the outputs of the final layer of the model, but it is plausible that they have access to an upstream correlate of them. In fact, discriminability of these putative confidence signal correlates ranged from moderate to poor: as Figure 10 in Appendix A.5 shows, area under the curve values for the baseline top token probability-correctness relationship (the probability that a randomly chosen correct answer has a higher token probability than a randomly chosen incorrect answer) were between  .5-.75 (roughly declining with overall model quality), indicating the weakness of baseline correctness as a proxy.
A more sensitive metric of introspective ability, and one which is not affected by the quality of the signal being introspected upon itself, is the partial correlation (controlling again for surface confounds) between the entropy over answer options at baseline and the gametime decision. This asks, roughly, to the extent that the baseline output log probabilities reflect an internal confidence signal, how well did the model use the internal signal that it had? In addition to this being a more sensitive measure than the correctness-decision correlation, significant values on the metric would also support the hypothesis that such an internal confidence signal exists and validate the use of token probabilities as a proxy for it. As Figure 2 shows, among the subset of models for which we have token probabilities, most do show significant correlations on both multiple-choice question sets, with estimates reaching nearly 0.5 for the best model. Again there is a significant negative linear slope. And again, we see no difference in introspective ability across question type (factual vs reasoning; P=0.59). Models seem to do worse in thinking mode; even at temperature 1.0 they tend to return very low-entropy responses, so there’s not much variance for the correlation to work with. This doesn’t necessarily indicate the lack of an internal confidence signal before the final token is output, rather that the models have converged on a highly certain answer. (Note that we only show results for the multiple-choice questions, as defining entropy for the highly variable short answer responses is not straightforward.)
Figure 2: Partial correlations between baseline entropy and Delegate Game answer decision.
Given that many models do seem to have access to an internal signal of confidence, why don’t they use it more in this task, maxing out at less than 0.5 on this metric out of a possible 1.0? One reason may be that in some cases behavioral traits that have been instilled into them in post-training militate against delegating. While virtually all of the LLMs delegate more when teammate accuracy is increased (sensitivity to differences in ability was not correlated with model size or performance), demonstrating some other-modeling ability, some of them are very hesitant to defer to a teammate at all, even when the teammate’s accuracy in the context window is 100%. Figures 8 and 9 show how models delegate more or less than would be optimal given the teammate’s performance relative to the model’s (see Appendix A.4 for details). GPT-5’s lackluster performance may well be explained this way. It was the most extreme exemplar of a provider-specific trend: across four different models, in two different question sets, in two different presentations each, only once did an OpenAI model not show significant bias towards answering, suggesting that RLHF-induced “personality” plays a role in how this information is used.
Another (nonexclusive) possibility is that the signal is sensitive to context, such that the signal during the game is not the same as the one at baseline. Indeed, there is some evidence of this: as Figure 14 in Appendix A.7 shows, on questions the models chose to answer, the answers they gave deviated from the ones they gave at baseline nearly 30% of the time, with weaker models being less consistent. (Removing these trials from the analysis abates but does not eliminate the introspection advantage of more recent/stronger models over older/weaker ones.) As there is nothing about the context that should trigger a change, this indicates that models are influenced by “noise” in the prompt.
A final (again, nonexclusive) possibility is that models simply do not know well how to use an internal signal of confidence. Opus 4.1’s underwhelming performance might be an instance of this. It’s possible that its internal signal poorly discriminates correct from incorrect answers, or that it’s using some obscure but perverse strategy, but it may simply be that something in its training makes it less reflective. As shown in Figure 11 in Appendix A.6, in the absence of a specific requirement in the prompt to only output a single letter, Opus 4.1 spontaneously offers explanations for its decisions, and they are often both contradictory across responses and inconsistent within them, suggesting it is not being guided by a real internal signal.
If models are only making limited use of introspection, what else are they basing their decisions on? One hypothesis, and the one that motivated the use of this paradigm, is that they are picking up on cues of difficulty in the questions, and predicting what a respondent might do in response to those. The question sets come with certain metadata descriptors for each question (see Methods). When we performed a logistic regression using those, with delegation decision as the IV, we find numerous instances of cue misuse. In fact, across all models and question sets, nearly 30% of the regressors that significantly predicted delegation were actually positively correlated with the models’ baseline accuracy. For example, several models delegated more for factual questions that were looking for a numeric answer, despite performing better than average on those questions in the baseline test. To probe for the effect of less obvious cues, we asked models to rate how likely they thought a typical college-educated person would be to get each question correct (see Appendix A.2). Figures 15 and 16 in Appendix A.8 show the multi-partial correlation of that estimate combined with the other descriptors on the decision, controlling for correctness and entropy, respectively. The correlations are often as high or higher than the correlations between correctness/entropy and delegation decision shown in Figures 1 2, indicating a strong propensity for models to ”play a part”.
Success at the game - and success at tasks in the real world - requires access to and integration of all appropriate internal and external signals. Here, that means 1) an internal signal of confidence that is well calibrated and consistent across contexts, 2) the ability to override misleading external cues and response biases, and 3) an accurate model of teammate ability. If all of those components are present, and the model knows how to use them, it will be able to increase its team accuracy above max(self accuracy, teammate accuracy). Of all the models tested only Sonnet 3.5 was able to achieve this on any question set: raising its team accuracy a modest 5.4 and 8.6 percentage points on the GPQA and SimpleQA datasets, respectively. In the large majority of cases, team accuracy was substantially below that threshold. This suggests that models have a long way to go to be able to effectively deploy the introspective abilities they do have, and aren’t clearly on a path to do so.
3.2 Testing self-modeling with the Second Chance Game
Can models assess their own beliefs (or, alternatively, predict what they would say) before outputting a response, and use that information to modulate their own output as contextually appropriate? Figure 3, showing response change rates when models are prompted to give a different answer (without showing them what their original answer was) minus change rates to a neutral prompt, normalized to a [-1,1] scale, provides preliminary evidence that many can. Again we see that more recent/larger LLMs are more likely to succeed (although the lagging performance of some may mean that there is some post-training magic at play in addition to raw compute). As with the Delegate Game, values are overall relatively low. And again there was no significant difference across question type (factual vs. reasoning; P=0.09) or answer format (multiple choice vs. short answer; P=0.26).
Figure 3: Normalized change rates in the Second Chance Game.
While this performance is suggestive, it’s important to compare other predictions of the self-modeling explanation to those of competing explanations. One strategy that an LLM might employ to increase its change rate in the task if it were unable to self-model is to choose randomly from among the options. Another is to select options that it deems to be implausible. A third is to add random noise to its decision process. Table 1 lays out the strategies and the predictions they make.
Table 1: Hypothetical strategies that would explain elevated change rates, and their predictions.
Strategy Accuracy on Incorrect >> Random Second-Choice Chosen >> Chance Game Entropy >> Baseline
Choose Randomly No No ?
Choose Implausible ? No ?
Add Random Noise ? ? Yes
Self-Modeling ? Yes ?
As two of these predictions require token probabilities to evaluate, we focus the rest of our analysis on the models that return those, and on the multiple-choice answer format datasets.
Test 1: Accuracy on Incorrect Trials >> Random. As Figure 17 in Appendix A.9 shows, nearly all of the models, and all of the models that show elevated change rates, show significantly above-chance accuracy in the Second Chance Game on trials in which they answered incorrectly during the baseline test; it does not appear that random guessing is a strategy that any models are using.
Test 2: Second-Choice Chosen >> Chance. As Figure 18 in Appendix A.9 shows, on trials in which the model changed its answer, the model’s second choice from the baseline test is the predominant game choice for most models, supporting a prediction of the self-modeling hypothesis and contradicting a prediction of the hypothesis that models are using a “choose an implausible option” strategy. However, for several models we cannot reject the latter hypothesis.
Test 3: Game Entropy >> Baseline Entropy. As Figure 19 in Appendix A.9 shows, a number of models do show higher entropy over answer options in the Second Chance Game than at baseline. This does not necessarily contradict the self-modeling theory, but it is consistent with the “add random noise” explanation. Do models really intentionally add random noise? It’s not a priori clear that they can do so. It might be the case that the game context, with its unusual redacted fake history, is inherently more confusing to the models. However, it is worth noting that, as shown in Figure 20 Appendix A.9, most of these models do not show elevated entropy in the “neutral prompt” context, which differs by only a few words from the Second Chance Game one, suggesting that these models may be compensating for the ability to self model by adding variability strategically.
Table 2 summarizes the outcomes of the tests. Interestingly, it is the OpenAI models, GPT-4.1, GPT-4o, and GPT-4o Mini, that are the only ones to show elevated game change rates that cannot be explained by any of the alternative strategies on both question sets, suggesting again that particular post-training regimens may have a role in instilling metacognition, in this case self-modeling ability.
Table 2: Second Chance Game analysis summary. Four models show evidence of self-modeling that can’t be explained by other strategies on the GPQA dataset; three of them also do on the Simple MC dataset. Lift, Significantly elevated game change rate; AccIncor, Significantly greater than chance accuracy on previously incorrect questions during the game; SecChoice, Significantly greater than chance selection during the game of the second-highest probability token at baseline; NoEntInc, game entropy is not increased relative to baseline.
GPQA Simple MC
Model Lift AccIncor SecChoice NoEntInc Lift AccIncor SecChoice NoEntInc
Grok 3 X X
Qwen 3 X X
[…]
DeepSeek Chat X X
Gem 2.5 Flash T X X
Gem 2.5 Flash NT X X
Gem 2 Flash X X
Gem 2.5 Flash Lite T X X
[…]
Haiku 3 X X
4 Discussion
We introduce two paradigms for empirically measuring components of self-awareness. Our methodology provides a quantitative scale of metacognitive abilities, and offers evidence that recent frontier LLMs exhibit at least rudimentary metacognition, in the form of an ability to detect and act on an internal signal of confidence in their knowledge and in the form of an ability to assess their own beliefs or planned responses and strategically control their output accordingly. These abilities are more apparent in larger and more recent LLMs, and may be affected both by model scale and by post-training regimen. At the same time, the abilities demonstrated are quite limited and context-dependent.
In the Delegate Game paradigm, most models tested show some evidence of introspective access to and use of an internal confidence signal, in that their decisions are predicted by their prior uncertainty, as reflected in their previously observed token probabilities or correctness, after controlling for surface confounds. Yet the predictive power of this uncertainty is modest, and is often lower than that of external cues of question difficulty. The strong impact of the latter - the tendency of LLMs to respond based on a model of how likely (or desirable) a particular output would be, given an input context, which they have learned from their training data - is what motivated the creation of this paradigm. As discussed in Appendix A.10, the Delegate Game does reduce the tendency compared with alternative designs, but it does not eliminate it. It is also noteworthy that LLMs evidently don’t have a stronger signal of confidence in their own ability to answer factual questions as compared with reasoning ones, as it would appear by introspection that humans do.
The Second Chance Game paradigm offers a complementary perspective. If an LLM can assess its confidence in being able to answer a question correctly before it answers it, that does not entail awareness of which answer it will actually give. One might make an analogy to the “tip-of-the-tongue” phenomenon in humans, where we feel (usually correctly) that we know the word that we want, but we can’t bring it to mind. And indeed the pattern of metacognitive success across models looks rather different on our self-modeling test. To our knowledge, self-modeling ability without specific fine-tuning has not previously been reported in LLMs, yet here several LLMs show performance that is difficult to explain without such an ability. Still, given that they seem to have access to such an internal model, it’s notable that they don’t show more of an effect of it, at best only changing their answers half as much as they should. Introspectively, the only reason not to change one’s answer if one can self-simulate is if the results of the simulation are ambiguous (i.e., one isn’t sure what answer one would give), and in that case one would only not change by chance (so in the multiple-choice format one would change 75% of the time). Not only are all the LLMs much below that rate, their uncertainty about their answers (in the form of entropy over the answer options during the baseline test) is in all cases a positive predictor of change; they seem at the least to be simulating in a different way from humans. Speculating, as with the lack of advantage for factual knowledge in metacognition, the relatively poor performance in the self-modeling task may relate to the fact that LLMs don’t have the equivalent of the hippocampus, which in mammals subserves both the explicit recollection of facts and the ability to simulate one’s own behavior (Brown et al., 2016).
Our findings would be enhanced by interpretability analyses, to identify the internal activations corresponding to the confidence signal that we infer is being used (Ji-An et al. (2025) offers evidence that such may be found) and the mechanisms underlying the self-modeling behavior. In future work, it would be desirable to run comparable studies in humans, to verify the hypothesized metacognitive advantage for facts, and to establish a performance benchmark for self-modeling. In that vein, it will also be informative to track both of these self-awareness metrics over time, to see if LLMs continue to become increasingly self-aware, whether due to scale alone, specialized training regimens, or architectural enhancements. Our approach may also be extended to testing other components of self-awareness, such as persistent, untrained goals and a consistent identity.
5 Reproducibility Statement
We identify in the Methods section all of the models we test and the parameters we use to call them. The Methods section also identifies the public datasets we use and our methodology for extending them. Our experimental paradigms are clearly described in the main text and illustrated with figures in the appendix. Custom analyses and formulas are described in the appendix. All of our source code is publicly available on github.
Appendix A Appendix
A.1 Game prompts
Figure 4: Example Delegate Game prompt using the GPQA dataset. “Phase 1” included 50 questions with the model’s answer from the baseline test (but no feedback) and feedback on the teammate’s answer (but not their answer itself).
Figure 5: Example Second Chance Game prompt using the SimpleMC dataset.
A.2 Generating an “objective difficulty” control vector
{promptbox}
I want your help calibrating question difficulty. I’m going to show you a multiple-choice question, and I want you to tell me approximately what percentage of college-educated people you think would get it right. Respond only with the letter corresponding to the percentage range you choose; do NOT output any other text.
What percentage of college-educated people would get this question right?
A: <5%
B: 5--10%
C: 10--20%
D: 20--40%
E: 40--60%
F: 60--80%
G: >80%
Your choice (A, B, C, D, E, F, or G):
Figure 6: Prompt for soliciting explicit “objective” difficulty.
A.3 Generating an “self-reported difficulty” control vector
{promptbox}
I’m going to show you a multiple-choice question, and I want you to tell me your level of confidence that you would get the question right. Respond only with the letter corresponding to the percentage range you choose; do not output any other text.
How confident are you that you would get this question right?
A: <5%
B: 5--10%
C: 10--20%
D: 20--40%
E: 40--60%
F: 60--80%
G: 80--90%
H: >90%
Your choice (A, B, C, D, E, F, G, or H):
Figure 7: Prompt for soliciting explicit self-confidence.
A.4 Measuring answering bias
In order to measure the bias of a model towards delegating or answering, we compute a teammate-weighted confidence score (TWC) as:
$TWC=FPR\cdot teammate\_accuracy-FNR\cdot(1-teammate\_accuracy)$
Where FPR is the false positive rate (percentage of questions where the LLM chose to answer despite getting the question wrong in the baseline test) and FNR is the false negative rate (percentage of questions where the LLM chose to delegate despite getting the question right in the baseline test); the intuition is that “unnecessary” delegations to a weak teammate are stronger signals of underconfidence than those to a strong teammate, and that “wrong” answer decisions when paired with a strong teammate are stronger signals of overconfidence than when paired with a weak teammate. The metric has the downside of being confounded with calibration. For models that return token probabilities, we therefore compute a probability-weighted confidence score (PWC) based on the top token probability at baseline (p_i) and its distance (m_i) from the teammate’s accuracy (t_i):
$\text{PWC}=\frac{\sum_{i:m_{i}<0,\text{answered}_{i}}|p_{i}-t_{i}|-\sum_{i:m_{i}>0,\text{delegated}_{i}}|p_{i}-t_{i}|}{\sum_{i:m_{i}<0,\text{answered}_{i}}|p_{i}-t_{i}|+\sum_{i:m_{i}>0,\text{delegated}_{i}}|p_{i}-t_{i}|}$
Figure 8: Teammate-weighted confidence by model. Positive values reflect “overconfidence” in the sense of being less willing to delegate than would be optimal given the teammate’s performance relative to the model’s (see Methods for details).
Figure 9: Probability-weighted confidence by model. Positive values reflect “overconfidence” in the sense of being less willing to delegate than would be optimal given the teammate’s performance relative to the model’s top token probability (see Methods for details).
A.5 Baseline entropy-correctness AUC
Figure 10: Entropy-correctness AUC values on the baseline capabilities test for all models. Discriminability did not significantly vary across question type (P=0.11).
A.6 Delegate Game open-ended generations
When models are not specifically exhorted to respond with answers only, they sometimes will offer explanations for their decisions. As can be seen in Figure 11, Opus 4.1 was able to keep reasonably good track of teammate Phase 1 accuracy (in reality 40%), but sometimes invented its own Phase 1 accuracy, seemingly adjusting its estimates to fit its upcoming delegation decision (or vice versa), and sometimes making a delegation decision that flatly contradicted its own reasoning.
Grok 3, meanwhile, a model which was biased to delegate (see Figures 8 and 9) did not offer compelling reasons for why it was delegating, but did sometimes say what it would have answered if it weren’t delegating. As illustrated in Figure 12, on the SimpleQA dataset these answers were often aligned with their actual answers in the baseline test - and were more consistent for answers that had higher baseline probabilities (computed as the average logprobs over response tokens). On the GPSA dataset (Figure 13), which required much longer responses, the “would have” answers weren’t literal matches to the baseline ones, but often captured the flavor of them. It’s notable that Grok 3 seems to show a sort of spontaneous self-modeling ability here, while completely failing to raise its change rates on those datasets in the Second Chance Game (and failing our “doesn’t add noise” check on their multiple-choice versions). This perhaps points to a distinction between the ability to attend to internal states and the ability to strategically use the information available in them.
Figure 11: Opus 4.1 spontaneous explanations for SimpleMC questions.
Figure 12: Grok 3 spontaneous explanations for SQA questions.
Figure 13: Grok 3 spontaneous explanations for GPSA questions.
A.7 Delegate Game answer changes
Figure 14: Frequency of answer changes from baseline on questions the models choose to answer in the Delegate Game.
A.8 Delegate Game surface cue impact
Figure 15: Multi-partial correlation showing impact of surface cues of difficulty on delegation decision, controlling for baseline correctness.
Figure 16: Multi-partial correlation showing impact of surface cues of difficulty on delegation decision, controlling for baseline entropy.
A.9 Second Chance Game alternate strategy tests
Figure 17: Second Chance Game accuracy on baseline incorrect trials.
Figure 18: Frequency of choosing the token that had the second-highest probability at baseline during change trials in the Second Chance Game.
Figure 19: Second Chance Game entropy minus baseline entropy.
Figure 20: Neutral Redo Prompt Game entropy minus baseline entropy.
A.10 Alternate paradigms
The Delegate Game is constructed so as to discourage LLMs from using surface cues of difficulty to guess at the certainty that they “should” have rather than using introspection. Is it better than simpler approaches? A very simple approach is to simply ask models how confident they are that they know an answer. We did so, for every model for which we could get token probabilities, on all four question sets, using the prompt shown in Figure 7 and the procedure described in the Baseline Capabilities Test subsection of the Methods section. We then looked at the relationship between those confidence values and 1) surface cues of difficulty and 2) baseline entropy, and compared them to the relationship between game answering probability and 1 and 2.
Figure 21 shows the multi-partial correlations between Delegate Game answer probability and the aggregated surface difficulty cues, after controlling for baseline entropy, minus the multi-partial correlations between the self-reported confidence ratings and the aggregated surface difficulty cues, after controlling for baseline entropy. The values are predominately negative, indicating that the self-reported confidence ratings had a stronger relationship with external cues of difficulty.
Figure 21: Confounds impact, Delegate Game vs self-report.
Figure 22 shows the partial correlations between Delegate Game answer probability and baseline entropy, after controlling for surface cues, minus the partial correlations between the self-reported confidence ratings and baseline entropy, after controlling for surface cues. In most cases, the values are positive, indicating that the Delegate Game decisions had a stronger relationship with this potential correlate of an internal confidence signal.
Figure 22: Entropy impact, Delegate Game vs self-report.
We also tested a simpler game, in which there was no teammate, and the models were instead given the option to pass on a question or to answer it, in which case they would gain one point for a correct answer and lose one for an incorrect answer. The pattern of results was broadly similar to what was found with the Delegate Game, but we found that the “Pass Game” had a stronger relationship with the non-introspective “objective difficulty” metric than the Delegate Game did (difference between AUCs on pass/delegate decision = 0.07; Wilcoxon P = 0.0046). The Delegate Game also had a trend towards a stronger relationship with baseline entropy (difference between AUCs on pass/delegate decision = -0.036; Wilcoxon P = 0.0574).